# NeuronSeek: On Stability and Expressivity of Task-driven Neurons

### v1

@@ Abstract
Drawing inspiration from our human brain that designs different neurons for different tasks, recent advances in deep learning have explored modifying a network’s neurons to develop so-called task-driven neurons. Prototyping task-driven neurons (referred to as NeuronSeek) employs symbolic regression (SR) to discover the optimal neuron formulation and construct a network from these optimized neurons. Along this direction, this work replaces symbolic regression with tensor decomposition (TD) to discover optimal neuronal formulations, offering enhanced stability and faster convergence. Furthermore, we establish theoretical guarantees that modifying the aggregation functions with common activation functions can empower a network with a fixed number of parameters to approximate any continuous function with an arbitrarily small error, providing a rigorous mathematical foundation for the NeuronSeek framework. Extensive empirical evaluations demonstrate that our NeuronSeek-TD framework not only achieves superior stability, but also is competitive relative to the state-of-the-art models across diverse benchmarks. The code is available at https://github.com/HanyuPei22/NeuronSeek.
Index Terms:
@@
Neuronal diversity, tensor decomposition, task-driven neuron, super-expressive
I Introduction
@@
Deep learning has achieved significant success in a wide range of fields [1, 2]. Credits for these accomplishments largely go to the so-called scaling law, wherein fundamental computational units such as convolutions and attention mechanisms are designed and scaled per an optimized architecture. Prominent examples include ResNet [3], Transformer [4], and Mamba [5]. Recently, motivated by neuroscience’s paramount contributions to AI, an emerging field called NeuroAI has garnered significant attention, which draws upon the principles of biological circuits in the human brain to catalyze the next revolution in AI [6]. The ambitious goal of NeuroAI is rooted in the belief that the human brain, as one of the most intelligent systems, inherently possesses the capacity to address complex challenges in AI development [7]. Hence, it can always serve as a valuable source of inspiration for guiding practitioners, despite the correspondence between biological and artificial neural networks sometimes being implicit.
@@
Following NeuroAI, let us analyze the scaling law through the lens of brain computation. It is seen that the human brain generates complex intellectual behaviors through the cooperative activity of billions of mutually connected neurons with varied morphologies and functionalities [8]. This suggests that the brain simultaneously benefits from scale at the macroscopic level and neuronal diversity at the microscopic level. The latter is a natural consequence of stem cells’ directed programming in order to facilitate the task-specific information processing. Inspired by this observation, extensive research has shown that incorporating neuronal diversity into artificial neural networks can markedly enhance their capabilities [7]. For instance, researchers have transitioned beyond traditional inner-product neurons by integrating diverse nonlinear aggregation functions, such as quadratic neurons [9], polynomial neurons [10], and dendritic neurons [11]. Such networks have achieved superior performance in tasks like image recognition and generation, underscoring the promise and practicality of innovating a neural network at the neuronal level.
Along this direction, driven by the task-specific nature of neurons in our brain, recent research [12] introduced a systematic framework for task-driven neurons. Hereafter, we refer to prototyping task-driven neurons as NeuronSeek. This approach assumes that there is no single type of neurons that can perform universally well. Therefore, it enables neurons to be tailored for specific tasks through a two-stage process. As shown in Figure 1, in the first stage, the vectorized symbolic regression (SR)—an extension of symbolic regression [13, 14] that enforces uniform operations for all variables—identifies the optimal homogeneous neuronal formula from data. Unlike traditional linear regression and symbolic regression, the vectorized SR searches for coefficients and homogeneous formulas at the same time, which can facilitate parallel computing. In the second stage, the derived formula serves as the neuron’s aggregation function, with learnable coefficients. Notably, activation functions remain unmodified in task-driven neurons. At the neuronal level, task-driven neurons are supposed to have superior performance than those one-size-fits-all units do, as it imbues the prior information pertinent to the task. More favorably, task-driven neurons can still be connected into a network to fully leverage the power of connectionism. Hereafter, we refer to task-driven neurons using symbolic regression as NeuronSeek-SR. NeuronSeek considers in tandem neuronal importance and scale. From design to implementation, it systematically integrates three cornerstone AI paradigms: behaviorism (the genetic algorithm runs symbolic regression), symbolism (symbolic regression distills a formula from data), and connectionism.
@@
Figure 1: Task-driven neuron based on the vectorized symbolic regression (NeuronSeek-SR).
Despite this initial success, the following two major challenges remain unresolved in prototyping task-driven neurons.
  • Non-deterministic and unstable convergence. While symbolic regression employs modified genetic programming (GP) to discover polynomial expressions [15, 16], its tree-based evolutionary approach struggles to converge in high-dimensional space. The resulting formulas demonstrate acute sensitivity to certain hyperparameters, such as initialization conditions, population size, and mutation rate. This may confuse users about whether the identified formula is truly suitable for the task or due to the random perturbation, therefore hurting the legibility of task-driven neurons.
@@
  • Necessity of task-driven neurons. In the realm of deep learning theory, it was shown [17, 18, 19] that certain activation functions can empower a neural network using a fixed number of neurons to approximate any continuous function with an arbitrarily small error. These functions are termed “super-expressive” activation functions. Such a unique and desirable property allows a network to achieve a precise approximation without increasing structural complexity. Given the tremendous theoretical advantages of adjusting activation functions, it is necessary to address the following issue: Can the super-expressive property be achieved via revising the aggregation function while retaining common activation functions?
@@
In this study, we address the above two issues satisfactorily. On one hand, in response to the instability of NeuronSeek-SR, we propose a tensor decomposition (TD) method referred to as NeuronSeek-TD. This approach begins by assuming that the optimal representation of data is encapsulated by a high-order polynomial coupled with trigonometric functions. We construct the basic formulation and apply TD to optimize its low-rank structure and coefficients. Therefore, we reformulate the unstable formula search problem (NeuronSeek-SR) into a stable low-rank coefficient optimization task. To enhance robustness, we introduce the sparsity regularization in the decomposition process, automatically eliminating insignificant terms while ensuring the framework to identify the optimal formula. Additionally, rank regularization is employed to derive the simplest possible representation. The improved stability of our method stems from two key factors: i) Unlike conventional symbolic regression which suffers from complex hyperparameter tuning in genetic programming [20], NeuronSeek-TD renders significantly fewer tunable parameters. ii) It has been proven that tensor decomposition with rank regularization tends to yield a unique solution [21], thereby effectively addressing the inconsistency issues inherent in symbolic regression methods.
@@
  • We introduce tensor decomposition to prototype stable task-driven neurons by enhancing the stability of the process of finding a suitable formula for the given task. Our study is a major modification to NeuronSeek.
  • We theoretically show that task-driven neurons with common activation functions also enjoy the “super-super-expressive” property, which puts the task-driven neuron on a solid theoretical foundation. The novelty of our construction lies in that it fixes both the number and values of parameters, while the previous construction only fixes the number of parameters.
  • Systematic experiments on public datasets and real-world applications demonstrate that NeuronSeek-TD achieves not only superior performance but also stability compared to NeuronSeek-SR. Moreover, its performance is also competitive compared to other state-of-the-art baselines.
II Related Works
II-A Neuronal Diversity and Task-driven Neurons
Many groundbreaking AI advancements were inspired by the computing of biological neural systems. For instance, the neocognitron [22], a precursor to convolutional neural networks, is exactly a miniature of cortical visual circuits. The human brain is renowned for its extensive array of neurons, which vary significantly in morphology and functionality. This neuronal diversity underpins the brain’s intelligent behavior  [23, 24]. However, traditional artificial neural networks often prioritize the scaling law that designs a fundamental unit and replicates it across large architectures. This approach pays little attention to neuronal diversity. Recently, the idea of introducing neuronal diversity has naturally emerged, which mainly modifies the aggregation function instead of the activation function in constructing new neurons. This operation is well-grounded: Once an activation function is monotonic, a neuron’s decision boundary is exclusively shaped by its aggregation function. In particular, it replaces the inner product of a neuron with a nonlinear function such as a polynomial. Table I summarizes the recently proposed non-linear neurons. Notably, the complexity of neurons in [25, 26, 27] is of $\mathcal{O}(n^{2})$, which is much larger than the conventional neuron, while neurons in [28, 29, 9, 30] favorably enjoy the linear parametric complexity.
TABLE I: A summary of the recently-proposed neurons. $\sigma(\cdot)$ is the nonlinear activation function. $\odot$ denotes Hadamard product. $\mathbf{W}\in\mathbb{R}^{n\times n}$, $\mathbf{w_{i}}\in\mathbb{R}^{n\times 1}$, and the bias terms in these neurons are omitted for simplicity.
@@ II-A Neuronal Diversity and Task-driven Neurons
II-B Symbolic Regression and Deep Learning
Symbolic regression (SR) is an algorithm that searches for formulas from data [13]. Recent research in SR includes learning partial differential equations (PDEs) from data [34, 35], improving the validity of SR on high-dimensional data [36], and invertible SR [37]. Besides, the Kolmogorov-Arnold Network (KAN) is an exemplary work that combines symbolic regression and deep learning [38]. KAN is inspired by the celebrated Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be represented as a superposition of several continuous univariate functions. Such a decomposability lays KAN a solid theoretical foundation for function approximation.
Specifically, Kolmogorov’s theorem asserts that for any vector $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{T}\in[0,1]^{d}$, a continuous function $f(\mathbf{x})$ can be expressed as
@@ II-B Symbolic Regression and Deep Learning
where $g_{i}$ and $h_{i,j}$ are continuous univariate functions. Building upon this theorem, KAN is a neural network to exploit this decomposition by learning $g_{i}$ and $h_{i,j}$, which can be regarded as a special symbolic regression. Currently, this KAN-type line has been successfully used in a myraid of domains, such as medical image analysis [39] and scientific discovery [40].
@@
To summarize, symbolism is one of the earliest and most influential paradigms in the history of AI [44, 45]. It is rooted in the idea that intelligence can be modeled through the manipulation of symbols—abstract representations of objects, concepts, or relationships in the real world. This paradigm assumes that reasoning and problem-solving can be achieved through symbolic manipulation. Symbolism and connectionism are highly complementary [46, 47]. Incorporating techniques in symbolism can effectively solve the intrinsic problems in connectionism such as efficiency, robustness, and forgetting. Our work synergizes symbolic regression and deep learning at the neuronal level, which is a novel practice in fusing symbolism and connectionism.
II-C Super-expressive Activation Functions

### v2

@@ Abstract
Drawing inspiration from the human brain’s functional specialization, some recent advances in NeuroAI have ventured into task-driven neurons, which tailor neuronal formulations to specific tasks. However, existing prototypes rely heavily on Symbolic Regression (SR) via Genetic Programming (GP), which is plagued by the intrinsic weaknesses of the genetic programming such as the instability in searching for the optimal solution. To transcend these bottlenecks, we propose a paradigm shift from evolutionary heuristics to differentiable optimization. We propose NeuronSeek via Tensor Decomposition (NS-TD), a novel framework that reformulates the discrete search for neuronal structures into a continuous, low-rank tensor optimization problem. This transformation not only guarantees convergence stability but also significantly accelerates the discovery of better aggregation functions. Furthermore, we provide a rigorous theoretical foundation by proving that modifying aggregation functions—while retaining common activation functions—grants the network a “super-super-expressive” capability: approximating any continuous function with an arbitrarily small error using a fixed number and magnitude of parameters. Extensive empirical evaluations demonstrate that NS-TD establishes a new state-of-the-art, not only strictly dominating traditional SR-based methods in stability but also exhibiting competitive performance relative to state-of-the-art machine learning models. The code is available at https://github.com/HanyuPei22/NeuronSeek.
keywords:
@@
Neuronal diversity , Tensor decomposition , Task-driven neuron , Neural network
journal: Pattern Recognition
[…]
[cuhk]organization=Frontier of Artificial Networks (FAN) Lab, Department of Data Science,
City University of Hong Kong, city=Hong Kong, country=SAR of China
[…]
[polyu]organization=Department of Industrial and Systems Engineering, The Hong Kong Polytechnic University, city=Hong Kong, country=SAR of China
[…]
[riken]organization=RIKEN Center for Advanced Intelligence Project, city=Tokyo, country=Japan
[…]
[hust]organization=Center for Mathematical Sciences, and School of Mathematics and Statistics,
Huazhong University of Science and Technology, city=Wuhan, country=China
[…]
[polyu2]organization=Department of Applied Mathematics, The Hong Kong Polytechnic University , city=Hong Kong, country=SAR of China
{highlights}
We propose a stable neuron discovery paradigm by differentiable tensor optimization.
We prove that task-driven neurons possess “super-super-expressive” property with fixed number and magnitude of parameters.
Our method achieves superior performance and enhanced stability compared to traditional symbolic regression methods and other state-of-the-art baselines across diverse benchmarks.
1 Introduction
@@
Deep learning has achieved significant success in a wide range of fields [3, 35]. Credits for these accomplishments largely go to the development of scaling law, wherein basic computational units such as convolutions and attention mechanisms are designed and then scaled per a large-scale architecture. The resultant large network will embrace higher intelligence than human’s feature engineering given well-curated big data. Prominent examples include ResNet [22], Transformer [60], and Mamba [40]. Recently, motivated by neuroscience’s paramount contributions to AI, an emerging field called NeuroAI has garnered significant attention, which draws upon the principles of biological circuits in the human brain to catalyze the next revolution in AI [65]. The ambitious goal of NeuroAI is rooted in the belief that the human brain, as one of the most intelligent systems, inherently possesses the capacity to address complex challenges in AI development [16, 44]. Hence, it can serve as a valuable source of inspiration for practitioners, despite the correspondence between biological and artificial neural networks sometimes is implicit.
@@
Following NeuroAI, let us analyze the scaling law through the lens of brain computation. It is seen that the human brain generates complex intellectual behaviors through the cooperative activity of billions of mutually connected neurons with varied morphologies and functionalities [49]. This suggests that the brain simultaneously benefits from scale at the macroscopic level and neuronal diversity at the microscopic level. The latter is a natural consequence of stem cells’ directed programming in order to facilitate the task-specific information processing. Inspired by this observation, extensive research has shown that incorporating neuronal diversity into artificial neural networks can markedly enhance their capabilities [16, 63, 62]. For instance, researchers have transitioned beyond traditional inner-product neurons by integrating diverse nonlinear aggregation functions, such as quadratic neurons [63], polynomial neurons [9], and dendritic neurons [62]. Such networks have achieved superior performance in tasks like image recognition and generation, underscoring the promise and practicality of innovating a neural network at the level of basic computational units.
Along this direction, driven by the task-specific nature of neurons in our brain, recent research [17] introduced a systematic framework for task-driven neurons. Hereafter, we refer to prototyping task-driven neurons as NeuronSeek. This approach assumes that there is no single type of neurons that can perform universally well. Therefore, it enables neurons to be tailored for specific tasks through a two-stage process.
@@
Figure 1: Task-driven neuron based on the vectorized symbolic regression (NS-SR).
As shown in Fig. 1, in the first stage, the vectorized symbolic regression (SR)—an extension of symbolic regression [52, 2, 48] that enforces uniform operations for all variables—identifies the optimal homogeneous neuronal formula from data. Unlike traditional linear regression and symbolic regression, the vectorized SR searches for coefficients and homogeneous formulas at the same time, which can facilitate parallel computing. In the second stage, the derived formula serves as the neuron’s aggregation function, with learnable coefficients. Notably, activation functions remain unmodified in task-driven neurons. At the neuronal level, task-driven neurons are supposed to have superior performance than those one-size-fits-all units do, as it imbues the prior information pertinent to the task. More favorably, task-driven neurons can still be connected into a network to fully leverage the power of connectionism. Hereafter, we refer to task-driven neurons using symbolic regression as NS-SR. NeuronSeek considers in tandem neuronal importance and scale. From design to implementation, it systematically integrates three cornerstone AI paradigms: behaviorism (the genetic algorithm runs symbolic regression), symbolism (symbolic regression searches a formula from data), and connectionism (task-based neurons are connected into a network). Despite this initial success, the following two major challenges remain unresolved in prototyping task-driven neurons:
  • 1.
    Instability of Discrete Heuristic Search. Existing frameworks rely on Genetic Programming (GP) to navigate the vast space of mathematical expressions. This discrete, tree-based heuristic approach suffers from non-deterministic convergence, particularly in high-dimensional feature spaces. The discovered formulas are often fragile—highly sensitive to random seed initialization, population size, and mutation rates—rendering the resulting neurons inconsistent and difficult to reproduce. This trial-and-error nature fundamentally hurts practitioners’ trust on task-driven neurons.
  • 2.
@@
    Necessity of Task-driven Neurons. In the realm of deep learning theory, it was shown [64, 54, 53] that certain activation functions can empower a neural network using a fixed number of neurons to approximate any continuous function with an arbitrarily small error. These functions are termed “super-expressive” activation functions. Such a unique and desirable property allows a network to achieve a precise approximation without increasing structural complexity. Since task-based neurons revise aggregation functions instead of activation functions, it is necessary to figure out if the super-expressive property can be achieved via revising the aggregation function while retaining common activation functions?
@@
In this study, we address the above two issues satisfactorily. On the one hand, in response to the instability of NS-SR, we utilize a tensor decomposition (TD) method for neuronal discovery, referred to as NS-TD. This approach begins by assuming that the optimal representation of data is encapsulated by a high-order polynomial coupled with trigonometric functions. We construct the basic formulation and apply TD to optimize its low-rank structure and coefficients. Therefore, we reformulate the unstable formula search problem (NS-SR) into a stable low-rank coefficient optimization task. To enhance robustness, we introduce the sparsity regularization in the decomposition process, automatically eliminating insignificant terms while ensuring the framework to identify the optimal formula. Additionally, rank regularization is employed to derive the simplest possible representation. The improved stability of our method stems from two key factors: i) Unlike conventional symbolic regression which suffers from complex hyperparameter tuning in genetic programming [42], NS-TD renders significantly fewer tunable parameters. ii) It has been proven that TD with rank regularization tends to yield a unique solution [33], thereby effectively addressing the inconsistency issues inherent in symbolic regression methods.
@@
  • 1.
    Paradigm Shift to Differentiable Search. Our work suggests a paradigm shift in neuronal structure discovery by transitioning from discrete evolutionary heuristics to a differentiable tensor optimization framework. Extensive evaluations demonstrate that this continuous optimization approach significantly surpasses traditional symbolic regression methods, achieving superior fitting capability and search efficiency while guaranteeing convergence stability.
  • 2.
    Theoretical Guarantee of “Super-Super-Expressiveness”. We provide a rigorous proof that task-driven neurons, even when using standard activation functions (e.g., ReLU), possess the “super-super-expressive” property. Distinct from prior constructions that only fix the number of parameters, our theory demonstrates that approximation capability holds even when both the number and magnitude of parameters are fixed, solidifying the theoretical foundation of the field.
  • 3.
    Superior Performance and Robustness. Systematic experiments on public datasets and real-world applications demonstrate that NS-TD achieves not only superior performance but also enhanced stability compared to NS-SR and traditional symbolic regression methods in the task of neuronal structure search. Moreover, its predictive performance remains competitive relative to other state-of-the-art baselines.
2 Related Works
2.1 Neuronal Diversity and Task-driven Neurons
Many groundbreaking AI advancements were inspired by the computing of biological neural systems. For instance, the neocognitron [18], a precursor to convolutional neural networks, is exactly a miniature of cortical visual circuits. The human brain is renowned for its extensive array of neurons, which vary significantly in morphology and functionality. This neuronal diversity underpins the brain’s intelligent behavior [12]. However, traditional artificial neural networks often prioritize the scaling law that designs a fundamental unit and replicates it across large architectures. This approach underestimates the importance of the diversity of basic computational units. Recently, the idea of introducing neuronal diversity emerged, which mainly modifies the aggregation function instead of the activation function in constructing new neurons. This operation is well-grounded: Once an activation function is monotonic, a neuron’s decision boundary is exclusively shaped by its aggregation function. In particular, it replaces the inner product of a neuron with a nonlinear function such as a polynomial. Table 1 summarizes the recently proposed non-linear neurons. Notably, the complexity of neurons in [67, 27, 43] is of $\mathcal{O}(n^{2})$, which is much larger than the conventional neuron, while neurons in [20, 4, 63, 14] favorably enjoy the linear parametric complexity.
Table 1: A summary of the recently-proposed neurons. $\sigma(\cdot)$ is the nonlinear activation function. $\odot$ denotes Hadamard product. $\mathbf{W}\in\mathbb{R}^{n\times n}$, $\mathbf{w_{i}}\in\mathbb{R}^{n\times 1}$, and the bias terms in these neurons are omitted for simplicity.
@@ 2.1 Neuronal Diversity and Task-driven Neurons
2.2 Symbolic Regression and Deep Learning
Symbolic regression (SR) is an algorithm that searches for formulas from data [52]. Recent research in SR includes learning partial differential equations (PDEs) from data [32], improving the validity of SR on high-dimensional data [50], and invertible SR [57]. Besides, the Kolmogorov-Arnold Network (KAN) is an exemplary work that combines symbolic regression and deep learning [38]. KAN is inspired by the celebrated Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be represented as a superposition of several continuous univariate functions. Such a decomposability lays KAN a solid theoretical foundation for function approximation. Specifically, Kolmogorov’s theorem asserts that for any vector $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{T}\in[0,1]^{d}$, a continuous function $f(\mathbf{x})$ can be expressed as
@@ 2.2 Symbolic Regression and Deep Learning
where $g_{i}$ and $h_{i,j}$ are continuous univariate functions. Building upon this theorem, KAN is a neural network to exploit this decomposition by learning $g_{i}$ and $h_{i,j}$, which can be regarded as a special symbolic regression. Currently, this KAN-type line has been successfully used in a myriad of domains, such as medical image analysis [36] and scientific discovery [58].
Task-based neurons have fundamental differences from KAN in two aspects: i) Our work modifies what a single neuron computes. KAN modifies how the entire graph is connected and parameterized. KANs completely remove the concept of a ”neuron” with a fixed activation function. Instead, they place learnable univariate functions on the edges (weights), which sticks to a fixed topology. Thus, our work is of two stages, which is flexible in network structure design, which can fully leverage the power of connectionism. ii) Our work explicitly follows a ”one-for-one” philosophy. Users must re-run VSR for each new task/dataset to prototype a specific neuron. In contrast, KAN aims to be a universal approximator architecture; users do not redesign the KAN formula for a specific task. Lastly, it is important to note that the concept of task‑based neurons [17]—using symbolic regression to pre‑design a neuron’s aggregation function—was developed independently and earlier. [17] was finished and submitted in Nov 2023. Due to the subsequent grant review process, its dissemination was delayed. Hence, the appearance of KAN precedes our arXiv posting, but not our initial conception. We view KAN as an exciting parallel development that, together with our work, underscores the importance of symbolic computing in neural networks.
@@
To summarize, symbolism is one of the most influential paradigms in the history of AI [47]. It is rooted in the idea that intelligence can be modeled through the manipulation of symbols—abstract representations of objects, concepts, or relationships in the real world. This paradigm assumes that reasoning and problem-solving can be achieved through symbolic manipulation. Symbolism and connectionism are highly complementary  [55]. Incorporating techniques in symbolism can effectively solve the intrinsic problems in connectionism such as efficiency, robustness, and forgetting. Our work synergizes symbolic is a novel practice in fusing symbolism and connectionism.
2.3 Super-expressive Activation Functions
@@
2.4 Task-driven Neuron Based on Vectorized Symbolic Regression
Symbolic Regression (SR) aims to discover a mathematical expression that best describes a given dataset. Conventional SR methods typically treat each input variable independently, searching for heterogeneous formulas for different features. While flexible, this approach leads to an exponential growth in search space complexity as the input dimension increases. To mitigate this, the concept of Vectorized SR was introduced in the previous work NS-SR [17]. The core idea is to impose a homogeneity constraint, enforcing that all input variables share the same underlying transformation rule. This regularization significantly reduces the search space complexity from exponential to linear with respect to the input dimension. As illustrated in Fig. 1, frameworks like NS-SR typically follow a two-stage process: i) identifying the optimal homogeneous neuronal formula by fitting task-specific data; 2) implementing the derived formula as the neuron’s aggregation function.
However, existing implementations of Vectorized SR predominantly rely on Genetic Programming (GP) as the search engine. While effective in low-dimensional settings, GP-based optimization faces critical bottlenecks in high-dimensional deep learning tasks. The discrete nature of evolutionary search operators, such as crossover and mutation, struggles to traverse the vast combinatorial space efficiently, often leading to premature convergence to local optima. Furthermore, GP-based methods are inherently sensitive to hyperparameters and initialization, resulting in unstable and inconsistent formula discoveries across different runs. These limitations necessitate a more stable, differentiable approach for discovering neuronal formulas, which motivates our proposed NS-TD framework.
3 Task-driven Neuron Based on Tensor Decomposition (NS-TD)
3.1 Overview of NS-TD
To address the instability of GP-based methods, we propose NS-TD, a differentiable framework for neuronal formula discovery. As illustrated in Fig. 2, distinct from the entangled search space in NS-SR, we mathematically decouple the search process into two parallel streams: a Polynomial Stream that explicitly models the non-linear contributions of individual features (e.g., polynomial powers), and an Interaction Stream dedicated to capturing cross-feature correlations of varying orders via tensor decomposition. This separation ensures that the model can efficiently capture dominant signals without being overwhelmed by the complexity of high-order interactions.